# Using Enriched Category Theory to Construct the Nearest Neighbour Classification Algorithm

### v1

@@ 5 Conclusion
The nascent field of Category Theory for Machine Learning has been growing in recent years. As Category Theory is predominantly concerned with mathematical structure, there is a hope that such techniques can improve our understanding of how Machine Learning algorithms operate. Previous works have demonstrated that there is value in this avenue of research, but there is currently not enough experience to indicate the correct way to apply Category Theory to the understanding of Machine Learning algorithms. In particular, there has not previously been an application of Enriched Category Theory in Machine Learning. With the construction of the Nearest Neighbours Algorithm, using tools from Enriched Category Theory, there is now a stronger indication that this area can provide valuable insight. Furthermore, the strategies used for the representation of information and reasoning about the construction of machine learning algorithms in this format suggests that the enriched structure offers a potentially more intuitive framework than other categorical attempts.
The simplicity of constructing the Nearest Neighbours Algorithm in this framework does add credence to the sense that the algorithm itself is an exceedingly natural approach to extending classifications. With the formulation of the Extended Nearest Neighbours Algorithm, it becomes a tantalising area of future work to ask if this algorithm continues to provide sensible classifications in other bases of enrichment. This motivation is part of the underpinning interest mentioned in the introduction of this work. Is it the case that machine learning requires fundamentally new algorithms to tackle stranger and stranger problems. Or is it that when suitably abstracted, a handful of algorithms might prove to be sufficient for the majority of case and that the engineering challenge comes in choosing the correct base of enrichment.
Another interesting outcome of this work is to indicate that Enriched Category Theory is a framework of reasoning that should be of more interest to both Machine Learning Experts and Mathematicians. Often derided as a more abstract formulation of the exceedingly abstract field of Category Theory, it can be seen that certain basis of enrichment create enriched categories which are more practically useful than other categorical notions. Furthermore, it indicates that an understanding of the interaction between hom-objects, functors, and profunctors can provide useful insights into the structuring of information and the meaning behind those structures. Even if one does not find the rigorous application of the theory useful, the intuition may prove helpful.
Acknowledgments and Disclosure of Funding
This work was partly funded by the grant “Early detection of contact distress for enhanced performance monitoring and predictive inspection of machines” (EP/S005463/1) from the Engineering and Physical Sciences Research Council (EPSRC), UK, and Senseye.

### v2

@@ 5 Conclusion
The nascent field of Category Theory for Machine Learning has grown in recent years. As Category Theory is predominantly concerned with mathematical structure, there is hope that such techniques can improve our understanding of Machine Learning algorithms. Previous works have demonstrated that there is value in this avenue of research. However, there are insufficient examples. In response, this paper provides one of the first constructions of an existing algorithm in the language enriched category theory, where both its modelling behaviour and selection are presented within the language of category theory. The paper also demonstrates how such a construction might inform future analysis of machine learning algorithms with category theory.
The ability of category theory to generalise algorithm behaviour has been demonstrated in both the large and small scale. As no part of the construction relied directly on the use of $Cost$, this work has introduced an infinite family of algorithms for any choice of base of enrichment.
V-NNA $:X\nrightarrow Y$
V-NNA $(y,x):=V((\mathbf{1}_{NY}\circ F^{*})(y,x),\ (T_{*}\circ F^{*})(y,x))$
Because the algorithm was constructed in the category of Lawvere metric spaces, its classification labels can also be attributed with a metric, inducing the behaviour of soft and dependent classification labels.
The ability of category theory to make certain patterns and relationships more evident is also demonstrated. By representing the NNA as a profunctor, it was possible to demonstrate how the behaviour of the $k$-NNA is just another example of the NNA when projected into a space of tuples. It was also possible to highlight the pattern of the extension triangle as a consequence of the description of the data as two functors. This can be seen as a specific example of the algebraic description of finding relation from independent to dependent variables, which will hopefully become more apparent as more examples of algorithms are found.
Finally, by representing the NNA inside the category of Lawvere metric spaces, it becomes clearer how the environment of the algorithm affects its behaviour. As the internal language is expressible through distance, addition, truncated subtraction, infimum, and supremum, the NNA is only capable of interacting with concepts expressed in these terms. The relationship bewteen an algorithms behaviours and the category it exists inside of has interesting implications for future work in machine learning explainability.